# A Calibration Tool for Refractive Underwater Vision

### v1

@@ Abstract
Many underwater robotic applications relying on vision sensors require proper camera calibration, i.e. knowing the incoming light ray for each pixel in the image. While for the ideal pinhole camera model all viewing rays intersect in a single 3D point, underwater cameras suffer from - possibly multiple - refractions of light rays at the interfaces of water, glass and air. These changes of direction depend on the position and orientation of the camera inside the water-proof housing, as well as on the shape and properties of the optical window, the port, itself. In recent years explicit models for underwater vision behind common ports such as flat or dome port have been proposed, but the underwater community is still lacking a calibration tool which can determine port parameters through refractive calibration. With this work we provide the first open source implementation of an underwater refractive camera calibration toolbox. It allows end-to-end calibration of underwater vision systems, including camera, stereo and housing calibration for systems with dome or flat ports. The implementation is verified using rendered datasets and real-world experiments.
I Introduction
Camera calibration is a prerequisite for many robotics and computer vision applications. Relating image content to real world objects and measurements requires knowledge about the image formation characteristics of each utilized camera. The parameters defining this image formation, typically focal length, principal point and lens distortion parameters, are estimated during calibration. When acquiring images underwater, the light must travel between water, glass and air, media with different refractive indices, causing the light to be refracted at the interfaces. For certain use cases, e.g. keeping the camera at fixed working distances from objects, the refractive distortion can be absorbed by the lens distortion parameters and focal length of conventional camera models if the whole setup is calibrated underwater [1, 2]. In the general case, and in particular for settings where robots observe objects from very close and very far, the pinhole approximation can introduce significant systematic errors, since the amount of refractive distortion depends on scene distance [3, 4].
@@ I Introduction
An alternative approach for underwater cameras is to therefore not only model the camera, but also the housing protecting it. Theory for calibrating the two main setups, housings with a flat glass pane ’Flat Port’ and a dome shaped one ’Dome Port’ has been developed in e.g. [6, 7, 8] and [5, 9]. Several underwater applications use these explicit housing models i.e. in Structure-From-Motion [10, 11, 12, 13], Multi-View-Stereo 3D reconstruction [14] or visual odometry [15]. These approaches require knowledge about the housing parameters, which are hard to accurately measure, but have great influence on the image formation. Even methods aiming to compensate refraction such as [16] or those that use a simplified model [12] either need to know the flat port normal or assume perfect alignment. Similarly, dome port housings avoid refraction when the entrance pupil of the lens is perfectly centered inside the dome. In both cases it is useful to verify the alignment of housing and camera, either to model it explicitly or to assure that the assembly is correct. Housing calibration allows to precisely estimate the relevant properties. Despite these advantages, and despite the fact that several approaches have been proposed which explicitly model the housing, there are no available implementations or toolboxes for underwater camera calibration to the best of our knowledge. Our proposed application is meant as a service to the community and aims to fill this gap. We provide a cross platform, stand-alone tool which allows straight forward camera, stereo and housing calibration for users with varying technical background. The application is provided as an open-source implementation [^1] https://cau-git.rz.uni-kiel.de/inf-ag-koeser/calibmar.
@@
II Previous Work
For regular in-air calibration several tools exist. The MATLAB toolbox [18] and OpenCV library [19] are commonly known examples, the latter of which is also used in the ROS camera calibration package. Next to a few proprietary, commercial offerings some open-source specialized solutions are available as well. There are tools focusing on multi-camera systems [20, 21], combining camera and IMU calibration [20] for robotics applications or calibrating omnidirectional [22] and multi-spectral cameras [23]. The prevalence of these tools shows the importance of camera calibration for many fields. While, as mentioned in the previous section, various underwater vision approaches explicitly require the camera housings and theory for their calibration exists, no tools for this use case have been published yet, which is the motivation for our work that is outlined in the next sections.
@@ II Previous Work
III Application Structure and Features
@@ III-C Stereo Calibration
III-D Auto-Calibration Target
@@
Employing 3D calibration targets reduces the chance for these degeneracies but such targets are a lot more difficult to manufacture, as their 3D geometry has to be known precisely to define the correct world coordinate system. A third strategy is camera auto-calibration, where instead of relying on known calibration targets, the correspondences between general image features across multiple images are utilized. For an overview of auto-calibration see e.g. [26]. The approach can use arbitrary image material to calibrate a camera, which makes it very flexible, but also suffers from a few non-obvious pitfalls: Certain camera motion can lead to degenerate cases, detected features might be unreliable, scenes may contain non- rigid geometry (e.g. operator moving in the background) which must be filtered out or feature matches may not cover the whole image which can leave parts of the distortion parameters undetermined. When calibrating underwater, the requirement to move a camera around or through a feature rich scene can also be impractical.
For application with the tool, we propose a simple masking strategy to combine the advantages of auto-calibration with the control provided by a calibration target. Instead of simply using all features in a scene as detected by e.g. a SIFT feature detector [35], features which are not in a region of interest are masked out. The region of interest is defined by the presence of an ArUco marker [36] in the image. If a marker is detected its area is enlarged naively (i.e. in pixel space, ignoring perspective) around its center by a default or user specified factor, creating a mask for features to include. This simple definition allows the usage of virtually any object as a calibration target, as long at it is rigid, has unique, SIFT detectable features and contains an ArUco marker. A practical suggestion is to generate multiple random dot patterns [21] designed to generate many SIFT features, add an ArUco marker to the center and attach them to an object. Examples are shown in figure 4. Masked, matched features are then processed in a Structure from Motion (SfM) pipeline as implemented in COLMAP [25], with an extension for refraction from [33]. In addition to the sparse point cloud and camera poses, running SfM will also yield the auto-calibrated camera parameters. The presence of markers is additionally used to pre-calibrate the camera intrinsics. Each detected marker is interpreted as a individual view of four known 3D points in a planar coordinate system. These 2D-3D correspondences are then processed with the calibration method described in section III-A. Initializing the SfM algorithm with the obtained focal length estimate improves image registration and convergence.
@@ III-E Usability Features
Flat Port
Dome Port
Lighting
Camera
IV Validation
IV-A Simulated Images
Refracted images are rendered using the physically based renderer Mitsuba [38]. For camera calibration validation the rendered images are artificially distorted using the radial distortion model. For stereo calibration intrinsics are considered to be known. For ”underwater” renderings only the geometric effects of refraction are modeled, radiometric effects such as attenuation or scattering are ignored. Feature extraction algorithms for boards and SIFT include intensity normalizing steps which minimize their influence. Rendered images consist of sets for dome and flat port underwater as well as an ”in-air” set to validate lens distortion. Each of them are generated for checkerboard and auto-calibration targets individually. The dome port is simulated with 5 cm radius and 6 mm thickness, the flat port with 2 cm distance and 14 mm thickness. For both, $(\mu_{a},\mu_{g},\mu_{w})=\mathrm{(1,1.473,1.334)}$. The underlying camera has a field of view of 73°, which corresponds to a focal length of 1297.3655, with a resolution of 1920x1080 and a principal point in the image center. For the auto-calibration target the dome decentering is initialized with $(0,0,0)$, the flat port normal with $(0,0,1)$ and double the ground truth interface distance (the distance of interface to camera projection center is usually not known exactly). All estimated values with the respective simulation ground truth is given in table I. Note that the ground truth interface normal represents a rotation around the y-axis by 5°. For the dome port $E_{c}||$ denotes the dome port decentering estimation error in m. For the flat port $E_{r}||$ gives the error in interface distance estimation in m and $\epsilon$ the angular error for normal estimation in degrees. All rendered image sets consist of 25 images each.
TABLE I: Rendered dataset validation
@@ IV-A Simulated Images
IV-B Tank Images
Housing calibration is verified from real images. They are taken looking through the ports on a lab tank, as illustrated in figure 5. Because mechanically measuring the absolute projection center inside the dome or flat port as a ground truth reference is very challenging we instead opted for a ”differential” (relative motion) experiment. After acquiring a first set of images the camera was moved inside the dome or tilted in front of the flat port by a controlled movement. For the dome port the camera was shifted by 10 mm in between sets, in the flat port setup the camera was tilted by 5°. Table II shows the estimated values for each target and set, as well as the dome displacement error $E_{c}||$ in m and the interface normal angular error $\epsilon$ in degrees, as the deviation from the expected value for each target.
TABLE II: tank dataset validation
@@ IV-B Tank Images
IV-C Discussion
Both the checkerboard and auto calibration can closely recover camera intrinsics in the synthetic images. Bigger differences are only visible in the housing calibration sets. The checkerboard calibration estimations of dome port decentering are correct to millimeter precision, which roughly matches the accuracy reported in [5]. Interface distance (only known for the synthetic set) and tilt are well estimated using the checkerboard (with known absolute 3D coordinates) as well. The remaining error is similar to the ones reported in [7, 6, 39]. Note that for auto-calibration (unknown 3D target), absolute distances can not be obtained as the scene scale is unknown. Even a perfectly known interface thickness (absolute, in m) is - in practice - a too weak constraint to exactly determine the other parameters absolutely. This has already been noted in [39] and [5]. The refractive 3D reconstructions will often show scales which do not exactly match the interface thickness, in case no other absolute 3D information is present. This problem could be tackled by including scene scale constraints into the auto-calibration (i.e. some known properties of 3D target). One obvious source is the markers on the target, for which a user provided scale could be used to restrict the optimization.
V CONCLUSION

### v2

@@ Abstract
Many underwater applications rely on vision sensors and require proper camera calibration, i.e. knowing the incoming light ray for each pixel in the image. While for the ideal pinhole camera model all viewing rays intersect in a single 3D point, underwater cameras suffer from - possibly multiple - refractions of light rays at the interfaces of water, glass and air. These changes of direction depend on the position and orientation of the camera inside the water-proof housing, as well as on the shape and properties of the optical window, the port, itself. In recent years explicit models for underwater vision behind common ports such as flat or dome port have been proposed, but the underwater community is still lacking a calibration tool which can determine port parameters through refractive calibration. With this work we provide the first open source implementation of an underwater refractive camera calibration toolbox. It allows end-to-end calibration of underwater vision systems, including camera, stereo and housing calibration for systems with dome or flat ports. The implementation is verified using rendered datasets and real-world experiments.
1 Introduction
Camera calibration is a prerequisite for many computer vision applications. Relating image content to real world objects and measurements requires knowledge about the image formation characteristics of each utilized camera. The parameters defining this image formation, typically focal length, principal point and lens distortion parameters, are estimated during calibration. When acquiring images underwater, the light must travel between water, glass and air, media with different refractive indices, causing the light to be refracted at the interfaces. For certain use cases, e.g. keeping the camera at fixed working distances from objects, the refractive distortion can be absorbed by the lens distortion parameters and focal length of conventional camera models if the whole setup is calibrated underwater [1, 2]. In the general case, and in particular for settings where objects are observed from varying viewing distances, the pinhole approximation can introduce significant systematic errors, since the amount of geometric distortion depends on scene distance [3, 4, 5] and on the interface to projection center distance for flat ports [6, 7]. Additionally refraction can cause other image degradation effects such as chromatic aberration [8].
@@ 1 Introduction
An alternative approach for underwater cameras is to therefore not only model the camera, but also the housing protecting it. Theory for ray-tracing through refractive interfaces has been described [10, 11, 12] and solutions for calibrating the two main setups, housings with a flat glass pane ’Flat Port’ and a dome shaped one ’Dome Port’ have been developed in e.g. [13, 14, 15] and [16, 9, 17]. Several underwater applications use these explicit housing models i.e. in Structure-From-Motion [18, 19, 20, 21, 22], Multi-View-Stereo 3D reconstruction [23] or visual odometry [24]. These approaches require knowledge about the housing parameters, which are hard to accurately measure, but have great influence on the image formation. Even methods aiming to compensate refraction such as [6] or those that use a simplified model [21] either need to know the flat port normal or assume perfect alignment. Similarly, dome port housings avoid refraction when the entrance pupil of the lens is perfectly centered inside the dome. In both cases it is useful to verify the alignment of housing and camera, either to model it explicitly or to assure that the assembly is correct. Housing calibration allows to precisely estimate the relevant properties. Despite these advantages, and despite the fact that several approaches have been proposed which explicitly model the housing, there are no available implementations or toolboxes for underwater camera calibration to the best of our knowledge. Our proposed application is meant as a service to the community and aims to fill this gap. We provide a cross platform, stand-alone tool which allows straight forward camera, stereo and housing calibration for users with varying technical background. The application is provided as an open-source implementation [^1] https://cau-git.rz.uni-kiel.de/inf-ag-koeser/calibmar.
@@
2 Previous Work
For regular in-air calibration several tools exist. The MATLAB toolbox [26] and OpenCV library [27] are commonly known examples. Next to a few proprietary, commercial offerings some open-source specialized solutions are available as well. There are tools focusing on multi-camera systems [28, 29], combining camera and IMU calibration [28] for robotics applications or calibrating omnidirectional [30] and multi-spectral cameras [31]. The prevalence of these tools shows the importance of camera calibration for many fields. While, as mentioned in the previous section, various underwater vision approaches explicitly require the camera housings and theory for their calibration exists, no tools for this use case have been published yet, which is the motivation for our work that is outlined in the next sections.
@@ 2 Previous Work
3 Application Structure and Features
@@ 3.3 Stereo Calibration
3.4 Auto-Calibration Target
@@
Employing 3D calibration targets reduces the chance for these degeneracies but such targets are more difficult to manufacture, as their 3D geometry has to be known precisely to define the correct world coordinate system for classical calibration. A third strategy is camera auto-calibration, where instead of relying on known calibration targets, the correspondences between general image features across multiple images are utilized. For an overview of auto-calibration see e.g. [34]. The approach can use arbitrary image material to calibrate a camera, which makes it very flexible, but also suffers from a few non-obvious pitfalls: Certain camera motion can lead to degenerate cases, detected features might be unreliable, scenes may contain non- rigid geometry (e.g. operator moving in the background) which must be filtered out or feature matches may not cover the whole image which can leave parts of the distortion parameters undetermined. When calibrating underwater, the requirement to move a camera around or through a feature rich scene can also be impractical.
For application with the tool, we propose a simple masking strategy to combine the advantages of auto-calibration with the control provided by a calibration target. Instead of simply using all features in a scene as detected by e.g. a SIFT feature detector [42], features which are not in a region of interest are masked out. The region of interest is defined by the presence of an ArUco marker [43] in the image. If a marker is detected its area is enlarged naively (i.e. in pixel space, ignoring perspective) around its center by a default or user specified factor, creating a mask for features to include. This simple definition allows the usage of virtually any object as a calibration target, as long at it is rigid, has unique, SIFT detectable features and contains an ArUco marker. A practical suggestion is to generate multiple random dot patterns [29] designed to generate many SIFT features, add an ArUco marker to the center and attach them to an object. Examples are shown in figure 4. Masked, matched features are then processed in a Structure from Motion (SfM) pipeline as implemented in COLMAP [33], with an extension for refraction from [40]. In addition to the sparse point cloud and camera poses, running SfM will also yield the auto-calibrated camera parameters. The presence of markers is additionally used to pre-calibrate the camera intrinsics. Each detected marker is interpreted as a individual view of four known 3D points in a planar coordinate system. These 2D-3D correspondences are then processed with the calibration method described in section 3.1. Initializing the SfM algorithm with the obtained focal length estimate improves image registration and convergence.
@@ 3.5 Usability Features
4 Validation
Table 1: Rendered dataset validation
@@ 4 Validation
Table 2: Tank dataset validation
@@
4.1 Simulated Images
Refracted images are rendered using the physically based renderer Mitsuba [44]. For camera calibration validation the rendered images are artificially distorted using the radial distortion model. For stereo calibration intrinsics are considered to be known. For ”underwater” renderings only the geometric effects of refraction are modeled, radiometric effects such as attenuation or scattering are ignored. Feature extraction algorithms for boards and SIFT include intensity normalizing steps which minimize their influence. Rendered images consist of sets for dome and flat port underwater as well as an ”in-air” set to validate lens distortion. Each of them are generated for checkerboard and auto-calibration targets individually. The dome port is simulated with 5 cm radius and 6 mm thickness, the flat port with 2 cm distance and 14 mm thickness. For both, $(\mu_{a},\mu_{g},\mu_{w})=\mathrm{(1,1.473,1.334)}$. The underlying camera has a field of view of 73°, which corresponds to a focal length of 1297.3655, with a resolution of 1920x1080 and a principal point in the image center. For the auto-calibration target the dome decentering is initialized with $(0,0,0)$, the flat port normal with $(0,0,1)$ and double the ground truth interface distance (the distance of interface to camera projection center is usually not known exactly). All estimated values with the respective simulation ground truth is given in table 1. Note that the ground truth interface normal represents a rotation around the y-axis by 5°. For the dome port $E_{c}||$ denotes the dome port decentering estimation error in m. For the flat port $E_{r}||$ gives the error in interface distance estimation in m and $\epsilon$ the angular error for normal estimation in degrees. Stereo calibration is currently only implemented for the checker board target. All rendered image sets consist of 25 images each.
4.2 Tank Images
Flat Port
Dome Port
Lighting
Camera
Housing calibration is verified from real images. They are taken looking through the ports on a lab tank, as illustrated in figure 5. Because mechanically measuring the absolute projection center inside the dome or flat port as a ground truth reference is very challenging we instead opted for a ”differential” (relative motion) experiment. After acquiring a first set of images the camera was moved inside the dome or tilted in front of the flat port by a controlled movement. For the dome port the camera was shifted by 10 mm in between sets, in the flat port setup the camera was tilted by 5°. Table 2 shows the estimated values for each target and set, as well as the dome displacement error $E_{c}||$ in m and the interface normal angular error $\epsilon$ in degrees, as the deviation from the expected value for each target.
4.3 Discussion
Both the checkerboard and auto calibration can closely recover camera intrinsics in the synthetic images. Bigger differences are only visible in the housing calibration sets. The checkerboard calibration estimations of dome port decentering are correct to millimeter precision, which roughly matches the accuracy reported in [9]. Interface distance (only known for the synthetic set) and tilt are well estimated using the checkerboard (with known absolute 3D coordinates) as well. The remaining error is similar to the ones reported in [14, 13, 46]. Note that for auto-calibration (unknown 3D target), absolute distances can not be obtained as the scene scale is unknown. Even a perfectly known interface thickness is - in practice - a too weak constraint to exactly determine the other parameters absolutely. This has already been noted in [46] and [9]. The refractive 3D reconstructions will often show scales which do not exactly match the interface thickness, in case no other absolute 3D information is present. This problem could be tackled by including scene scale constraints into the auto-calibration (i.e. some known properties of 3D target). One obvious source is the markers on the target, for which a user provided scale could be used to restrict the optimization.
5 CONCLUSION